From past experience, Anna knows that the probability that her friend will take her to dinner in Lawiswis is 0.7, the probability that they will watch movies is 0.5, and the probability that he will take her to dinner or they will watch movies is 0.9. What is the probability that (1) they will do both? (2) they will do neither?
Answer by robertb(5830)


(1) 0.7 + 0.5 - 0.9 = 0.3
(2) 1 - 0.9 = 0.1